OpenStudy (anonymous):

medal If x varies directly with y and x = 3.5 when y = 14, find x when y = 18. 5 4 5.5 4.5

OpenStudy (anonymous):

can anyone help please so need this

OpenStudy (anonymous):

it's a ratio problem apply rule of three 3.5/x=14/18 =>x=(3.5*18)/14

OpenStudy (anonymous):

x is 4.5

OpenStudy (anonymous):

thank you thank you